OpenStudy (anonymous):

Veryify Gauss's theorem for the given three dimensional region D and vector field F. F= (x,y,z) D = {(x,y,z)|0<=z<=9-x^2-y^2}

OpenStudy (anonymous):

Gauss's theorem says: \[\int\limits_S \vec{V} \cdot d \vec{a} = \int\limits_V (\vec{\nabla} \cdot \vec{V}) dV\] Therefore you need to do both of these integrals.

OpenStudy (anonymous):

I am aware. I got the triple integral but I don't know how to set up the surface integral.

OpenStudy (anonymous):

Oh okay. Well you need to parametrize your surface with two independent parameters: \[\vec{\Phi}(x,y) = (x,y,9-x^2-y^2)\] From there you take the derivative with respect to the two parametrizations and cross them: \[\frac{\partial \vec{\Phi}}{\partial x} \times \frac{\partial \vec{\Phi}}{\partial y}\] Then evaluate your vector field at phi: \[\vec{F}(\vec{\Phi}(x,y))\] Then you have: \[\iint \vec{F}({\vec{\Phi}(x,y)}) \cdot \left( \frac{\partial \vec{\Phi}}{\partial x} \times \frac{\partial\vec{\Phi}}{\partial y}\right)dx dy\] Now to find the range on x and y you need to set z to zero giving: \[9 = x^2+y^2 \implies 0 \le x, y \le 3\]

OpenStudy (anonymous):

And don't forget to check whether it is orientation preserving or reversing! Though I don't really recall how to check :/

OpenStudy (anonymous):

Do you remember @zepdrix ? I know what an orientable surface is but that doesn't help with figuring out preserving/reversing...

zepdrix (zepdrix):

haha no i dunno XD I think this math is a smidgin above my current level

OpenStudy (anonymous):

From what I remember, the path needs to have the entire surface on the left side of the path which will show itself when I take the integral.

OpenStudy (anonymous):

Also doesn't my surface need to be closed so I would have to add a disk of radius 3 at the bottom?

OpenStudy (anonymous):

It is indeed, so you would need to add both parts. But for the other part its easy because you just have a parametrization of: \[\vec{\Phi}(r,\phi) = (r \cos(\phi), r \sin(\phi),0)\] As z = 0 for the disk.

OpenStudy (anonymous):

\[0 \le r \le 3; 0 \le \phi \le 2 \pi\]

OpenStudy (anonymous):

Yeah thats what I thought. Just wanted to be sure. Thank you

OpenStudy (anonymous):

No problem!